A reaction has an activation energy, Ea = 41.6 kJ mol-1 at 25 °C. By how many degrees must the temperature be raised in order to increase the reaction rate by a factor of 30.

  • A) 2.0
  • B) 374
  • C) 29
  • D) 76
  • E) 6.0
Answer #1

solution :- Activation energy, Ea = 41-6 kl/mol T= 25°C = 298k T2 - 2 y Kick then 12 = 30k using intigrated form af IC=A é Esа 12 = 373 - 70 % C - increase in temperature = T2-T = 573 - 70% - 29% = 7S. 10% - Нhsocc : 2 10